Financial statement data for years ending December 31 for Newton Company follow:

  20Y9 20Y8
Cash (end of year) $25,500 $24,250
Short-term investments (end of year) $8,270 $9,460
Operating expenses $60,135 $63,780
Depreciation expense $13,225 $11,400

Determine the days' cash on hand for 20Y8 and 20Y9. Round all calculations to one decimal place.
Answer #1

Calculate following

20Y9 20Y8
Operating expense-Depreciation/365 (60135-13225/365) = 128.5 (63780-11400)/365 = 143.5
Days cash on hand (25500+8270)/128.5 = 262.8 dats (24250+9460)/143.5 = 234.9 days